### v5

# Tree independence number III. Thetas, prisms and stars

Abstract.
We prove that for every $t\in\mathbb{N}$ there exists $\tau=\tau(t)\in\mathbb{N}$ such that every (theta, prism, $K_{1,t}$)-free graph has tree independence number at most $\tau$ (where we allow “prisms” to have one path of length zero).
§ Princeton University, Princeton, NJ, USA. Supported by NSF grant DMS-2348219 and by AFOSR grant FA9550-22-1-0083.
Department of Combinatorics and Optimization, University of Waterloo, Waterloo, Ontario, Canada.
Corresponding Author (shajebi@uwaterloo.ca).
CNRS, ENS de Lyon, Université Lyon 1, LIP UMR 5668, 69342 Lyon Cedex 07, France. Partially supported by the French National Research Agency under research grant ANR DIGRAPHS ANR-19-CE48-0013-01 and the LABEX MILYON (ANR-10-LABX-0070) of Université de Lyon, within the program Investissements d’Avenir (ANR-11-IDEX-0007) operated by the French National Research Agency (ANR), and H2020-MSCA-RISE project CoSP- GA No. 823748
1. Introduction
Graphs in this paper have finite and non-empty vertex sets, no loops and no parallel edges. The set of all positive integers is denoted by $\mathbb{N}$, and for every $n\in\mathbb{N}$, we write $[n]$ for the set of all positive integers no greater than $n$.
Let $G=(V(G),E(G))$ be a graph. A clique in $G$ is a set of pairwise adjacent vertices. A stable or independent set in $G$ is a set of vertices no two of which are adjacent. The maximum cardinality of a stable set is denoted by $\alpha(G)$, and the maximum cardinality of a clique in $G$ is denoted by $\omega(G)$. For a graph $H$ we say that $G$ contains $H$ if $H$ is isomorphic to an induced subgraph of $G$. We say that $G$ is $H$-free if $G$ does not contain $H$. For a set $\mathcal{H}$ of graphs, $G$ is $\mathcal{H}$-free if $G$ is $H$-free for every $H\in\mathcal{H}$. For a subset $X$ of $V(G)$, we denote by $G[X]$ the induced subgraph of $G$ with vertex set $X$, we often use “XX” to denote both the set $X$ of vertices and the graph $G[X]$.
Let $X\subseteq V(G)$. We write $N_{G}(X)$ for the set of all vertices in $G\setminus X$ with at least one neighbor in $X$, and we define $N_{G}[X]=N_{G}(X)\cup X$. When there is no danger of confusion, we omit the subscript “GG”. For $Y\subseteq V(G)$, we write $N_{Y}(X)=N_{G}(X)\cap Y$ and $N_{Y}[X]=N_{Y}(X)\cup X$. When $X=\{v\}$ is a singleton, we write $N_{Y}(v)$ for $N_{Y}(\{v\})$ and $N_{Y}[v]$ for $N_{Y}[\{v\}]$.
Let $x\in V(G)$ and let $Y\subseteq V(G)$. We say that $x$ is complete to $Y$ in $G$ if $N_{Y}[x]=Y$, and we say that $x$ is anticomplete to $Y$ in $G$ if $N_{G}[x]\cap Y=\emptyset$. In particular, if $x\in Y$, then $x$ is neither complete nor anticomplete to $Y$ in $G$. For subsets $X,Y$ of $V(G)$, we say that $X$ and $Y$ are complete in $G$ if every vertex in $X$ is complete to $Y$ in $G$, and we say that $X$ and $Y$ are anticomplete in $G$ if every vertex in $X$ is anticomplete to $Y$ in $G$. In particular, if $X$ and $Y$ are either complete or anticomplete in $G$, then $X\cap Y=\emptyset$.
For a graph $G=(V(G),E(G))$, a tree decomposition $(T,\beta)$ of $G$ consists of a tree $T$ and a map $\beta:V(T)\to 2^{V(G)}$ with the following properties:
  • For every $v\in V(G)$, there exists $t\in V(T)$ with $v\in\beta(t)$.
  • For every $v_{1}v_{2}\in E(G)$, there exists $t\in V(T)$ with $v_{1},v_{2}\in\beta(t)$.
  • $T[\{t\in V(T)\mid v\in\beta(t)\}]$ is connected for all $v\in V(G)$.
The treewidth of $G$, denoted $\operatorname{tw}(G)$, is the smallest integer $w\in\mathbb{N}$ such that $G$ admits a tree decomposition $(T,\beta)$ with $|\beta(t)|\leq w+1$ for all $t\in V(T)$. The tree independence number of $G$, denoted $\operatorname{{\mathsf{tree}-\alpha}}(G)$, is the smallest integer $s\in\mathbb{N}$ such that $G$ admits a tree decomposition $(T,\beta)$ with $\alpha(G[\beta(t)])\leq s$ for all $t\in V(T)$.
Both the treewidth and the tree independence number are of great interest in structural and algorithmic graph theory (see [ti1, tw15, ti2, dfgkm, dms2] for detailed discussions). They are also related quantitatively because, by Ramsey’s theorem [multiramsey], graphs of bounded clique number and bounded tree independence number have bounded treewidth (see also Lemma 3.2 in [dms2]). Dallard, Milanič, and Štorgel [dms2] conjectured that the converse is also true in hereditary classes of graphs (meaning classes which are closed under taking induced subgraphs). Let us say that a graph class $\mathcal{G}$ is $(\operatorname{tw},\omega)$-bounded if there is a function $f:\mathbb{N}\rightarrow\mathbb{N}$ such that every graph $G\in\mathcal{G}$ satisfies $\operatorname{tw}(G)\leq f(\omega(G))$.
Conjecture 1.1 (Dallard, Milanič, and Štorgel [dms2]).
For every hereditary class $\mathcal{G}$ which is $(\operatorname{tw},\omega)$-bounded, there exists $\tau=\tau(\mathcal{G})\in\mathbb{N}$ such that $\operatorname{{\mathsf{tree}-\alpha}}(G)\leq\tau$ for all $G\in\mathcal{G}$.
Conjecture 1.1 was recently refuted [lwnew] by two of the authors of this paper. It is still natural to ask: which $(\operatorname{tw},\omega)$-bounded hereditary classes have bounded tree independence number? So far, the list of hereditary classes known to be of bounded tree independence number is not very long (see [ti1, dms3, dms2] for a few). More hereditary classes are known to be $(\operatorname{tw},\omega)$-bounded. The reasons for the existence of the bound are often highly non-trivial, and it is not known whether the corresponding class has bounded tree independence number. A notable instance is the class of all (theta, prism)-free graphs excluding a fixed forest [tw8], which we will focus on in this paper.
Let us first give a few definitions. Let $P$ be a graph which is a path. Then we write, for $k\in\mathbb{N}$, $P=p_{1}\hbox{-}\cdots\hbox{-}p_{k}$ to mean $V(P)=\{p_{1},\ldots,p_{k}\}$, and for all $i,j\in[k]$, the vertices $p_{i}$ and $p_{j}$ are adjacent in $P$ if and only if |ij|=1|i-j|=1. We call the vertices $p_{1}$ and $p_{k}$ the ends of $P$, and we say that $P$ is a path from $p_{1}$ to $p_{k}$ or a path between $p_{1}$ and $p_{k}$. We refer to $V(P)\setminus\{p_{1},p_{k}\}$ as the interior of $P$ and denote it by $P^{*}$. The length of a path is its number of edges. Given a graph $G$, by a path in $G$ we mean an induced subgraph of $G$ which is a path. Similarly, for $t\in\mathbb{N}\setminus\{1,2\}$, given a $t$-vertex graph $C$ which is a cycle, we write $C=c_{1}\hbox{-}\cdots\hbox{-}c_{t}\hbox{-}c_{1}$ to mean $V(C)=\{c_{1},\ldots,c_{t}\}$, and for all $i,j\in[t]$, the vertices $c_{i}$ and $c_{j}$ are adjacent in $C$ if and only if $|i-j|\in\{1,t-1\}$. The length of a cycle is its number of edges (which is the same as its number of vertices). For a graph $G$, a hole in $G$ is an induced subgraph of $G$ which is a cycle of length at least four.
Figure 1. A theta (left) and a prism (right). Dashed lines represent paths of arbitrary (possibly zero) length.
A theta is a graph $\Theta$ consisting of two non-adjacent vertices $a,b$, called the ends of $\Theta$, and three pairwise internally disjoint paths $P_{1},P_{2},P_{3}$ of length at least two in $\Theta$ from $a$ to $b$, called the paths of $\Theta$, such that $P_{1}^{*},P_{2}^{*},P_{3}^{*}$ are pairwise anticomplete in $\Theta$ (see Figure 1). A prism is a graph $\Pi$ consisting of two triangles $\{a_{1},a_{2},a_{3}\},\{b_{1},b_{2},b_{3}\}$ called the triangles of $\Pi$, and three pairwise disjoint paths $P_{1},P_{2},P_{3}$ in $\Pi$, called the paths of $\Pi$, such that for each $i\in\{1,2,3\}$, $P_{i}$ has ends $a_{i},b_{i}$, for all distinct $i,j\in\{1,2,3\}$, $a_{i}a_{j}$ and $b_{i}b_{j}$ are the only edges of $\Pi$ with an end in $P_{i}$ and an end in $P_{j}$, and for every $i\neq j\in\{1,2,3\}$ $P_{i}\cup P_{j}$ is a hole (see Figure 1). If follows that if $P_{2}$ has length zero, then each of $P_{1},P_{3}$ has length at least two. We remark that the last condition is non-standard; the paths of a prism are usually of non-zero length, and a prism with a length-zero path is sometimes called a “line-wheel.” For a graph $G$, a theta in $G$ is an induced subgraph of $G$ which is a theta and a prism in $G$ is an induced subgraph of $G$ which is a prism.
The following was proved in [tw8] to show that the local structure of the so-called “layered wheels” [layered-wheels] is realized in all theta-free graphs of large treewidth. It also characterizes all forests, and remains true when only the usual “prisms” (with no length-zero path) are excluded:
Theorem 1.2 (Abrishami, Alecu, Chudnovsky, Hajebi, Spirkl [tw8]).
Let $F$ be a graph. Then the class of all (theta, prism, $F$)-free graphs is $(\operatorname{tw},\omega)$-bounded if and only if $F$ is a forest.
We propose the following strengthening (again, this may be true even with the usual “prisms” excluded):
Conjecture 1.3.
For every forest $F$, there is a constant $\tau=\tau(F)\in\mathbb{N}$ such that for every (theta, prism, $F$)-free graph $G$, we have $\operatorname{{\mathsf{tree}-\alpha}}(G)\leq\tau$.
As far as we know, Conjecture 1.3 remains open even for paths. But our main result settles the case of stars. For every $t\in\mathbb{N}$, let $\mathcal{C}_{t}$ be the class of all (theta, prism, $K_{1,t}$)-free graphs. We prove that:
Theorem 1.4.
[…]
2. Outline of the main proof
Like several earlier results [ti1, ti2, tw15] coauthored by the first two authors of this work, the proof of Theorem 1.4 deals with “balanced separators.” Let $G$ be a graph and let $w:V(G)\rightarrow\mathbb{R}^{\geq 0}$. For every $X\subseteq V(G)$, we write $w(X)=\sum_{v\in X}w(v)$. We say that that $w$ is a normal weight function on $G$ if $w(V(G))=1$. Given a graph $G$ and a weight function $w$ on $G$, a subset $X$ of $V(G)$ is called a $w$-balanced separator if for every component $D$ of $G\setminus X$, we have $w(D)\leq 1/2$. The main step in the proof of Theorem 1.4 is the following:
Theorem 2.1.
[…]
As shown below, Theorem 1.4 follows by combining Theorem 2.1 and the following (this is not a difficult result; see [ti2] for a proof):
Lemma 2.2 (Chudnovsky, Gartland, Hajebi, Lokshtanov and Spirkl; see Lemma 7.1 in [ti2]).
Let $s\in\mathbb{N}$ and let $G$ be a graph. If for every normal weight function $w$ on $G$, there is a $w$-balanced separator $X_{w}$ in $G$ with $\alpha(X_{w})\leq s$, then we have $\operatorname{{\mathsf{tree}-\alpha}}(G)\leq 5s$.
Proof of Theorem 1.4 assuming Theorem 2.1.
[…]
It remains to prove Theorem 2.1. The idea of the proof is the following. In [tw15] a technique was developed to prove that separators satisfying the conclusion of Theorem 2.1 exist. It consists of showing that the graph class in question satisfies two properties: being “amiable” and being “amicable.” Here we use the same technique. To prove that a graph class is amiable, one needs to analyze the structure of connected subgraphs containing neighbors of a given set of vertices. To prove that a graph is amicable, it is necessary to show that certain carefully chosen pairs of vertices can be separated by well-structured separators. Most of the remainder of the paper is devoted to these two tasks. Section 3 and Section 4 contain structural results asserting the existence of separators that will be used to establish amicability. Section 5 contains definitions and previously known results related to amiability. Section 6 contains the proof of the fact that the class $\mathcal{C}_{t}$ is amiable. Section 7 uses the results of Sections 3 and 4 to deduce that $\mathcal{C}_{t}$ is amicable, and to complete the proof of Theorem 2.1.
3. Breaking a wheel
A wheel in a graph $G$ is a pair $W=(H,c)$ where $H$ is a hole in $G$ and $c\in G\setminus H$ has at least three neighbors in $H$. We also use $W$ to denote the vertex set $H\cup\{c\}\subseteq V(G)$. A sector of the wheel $(H,c)$ is a path of non-zero length in $H$ whose ends are adjacent to $c$ and whose internal vertices are not. A wheel is special if it has exactly three sectors, one sector has length one and the other two (called the long sectors) have length at least two (see Figure 2 – A special wheel is sometimes referred to as a “short pyramid.”)
For a wheel $W=(H,c)$ in a graph $G$, we define the set $Z(W)\subseteq H\cup\{c\}$ as follows (see Figure 2). If $W$ is non-special, then $Z(W)=N_{H}[c]$. Now assume that $W$ is special. Let $ab$ be the sector of length one of $W$ and let $d$ be the neighbor of $c$ in $H\setminus\{a,b\}$. Then we define $Z(W)=\{a,b,c\}\cup N_{H}[d]$.
Figure 2. A non-special wheel $W$ (left) and a special wheel $W$ (right). Circled nodes represent the vertices in $Z(W)$.
Let $G$ be a graph. By a separation in $G$ we mean a triple $(L,M,R)$ of pairwise disjoint subsets of $V(G)$ with $L\cup M\cup R=V(G)$, such that neither $L$ nor $R$ is empty and $L$ and $R$ are anticomplete in $G$. Let $x,y\in V(G)$ be distinct. We say that a set $M\subseteq V(G)\setminus\{x,y\}$ separates $x$ and $y$ in $G$ if there exists a separation $(L,M,R)$ in $G$ with $x\in L$ and $y\in R$. Also, for disjoint sets $X,Y\subseteq V(G)$, we say that a set $M\subseteq V(G)\setminus(X\cup Y)$ separates $X$ and $Y$ if there exists a separation $(L,M,R)$ in $G$ with $X\subseteq L$ and $Y\subseteq R$. If $X=\{x\}$, we say that $M$ separates $x$ and $Y$ to mean $M$ separates $X$ and $Y$.
We have two results in this section; one for the non-special wheels and one for special wheels:
Theorem 3.1.
Let $G$ be a (theta, prism)-free graph, let $W=(H,c)$ be a non-special wheel in $G$ such that $H$ has length at least seven. Let $a,b\in G\setminus N[Z(W)]$ belong to (the interiors of) distinct sectors of $W$. Then $N[Z(W)]$ separates $a$ and $b$ in $G$.
Proof.
Let $S=N_{H}(c)$ and let $T=N[c]\cup(N[S]\setminus H)$. Then $T\subseteq N[Z(W)]$, and so it suffices to show that $T$ separates $a$ and $b$ (note that $a,b\notin T$). We begin with the following:
(1) Assume that some vertex $v\in G\setminus(W\cup T)$ has either a unique neighbor or two non-adjacent neighbors in some sector $P=p\hbox{-}\cdots\hbox{-}p^{\prime}$ of $W$. Let $b^{\prime}$ be the neighbor of $p$ in $W\setminus P$ and $b^{\prime\prime}$ be the neighbor of $p^{\prime}$ in $W\setminus P$. Then $N_{H}(v)\subseteq P\cup\{b^{\prime},b^{\prime\prime}\}$.
Otherwise, $v$ has a neighbor $d\in H\setminus(P\cup\{b^{\prime},b^{\prime\prime}\})$. Also, $c$ has a neighbor $d^{\prime}\in H\setminus(P\cup\{b^{\prime},b^{\prime\prime}\})$, as otherwise $W$ would be a prism or a special wheel. We choose $d$ and $d^{\prime}$ such that the path $Q$ in $H\setminus P$ from $d$ to $d^{\prime}$ is minimal. If $v$ has a unique neighbor $a$ in $P$, then $P\cup Q\cup\{c,v\}$ is a theta in $G$ with ends $a$ and $c$, a contradiction. Also, if $v$ has two non-adjacent neighbors in $P$, then $P\cup Q\cup\{v,c\}$ contains a theta with ends $c$ and $v$. This proves (3).
(2) For every $v\in G\setminus(W\cup T)$, there exists a sector $P$ of $W$ such that $N_{H}(v)\subseteq P$.
Suppose there exists a sector $P=p\hbox{-}\cdots\hbox{-}p^{\prime}$ such that $v$ has two non-adjacent neighbors in $P$. Then, by (3), we may assume up to symmetry that $v$ is adjacent to the neighbor $b$ of $p$ in $H\setminus P$. By (3), $b$ is the unique neighbor of $v$ in some sector $Q$ of $W$. So the fact that $v$ has at least two neighbors in $P$ contradicts (3) applied to $v$ and $Q$.
Suppose there exists a sector $P=p\hbox{-}\cdots\hbox{-}p^{\prime}$ such that $v$ has a unique neighbor $a$ in $P$. By (3), we may assume that $N_{H}(v)=\{a,b^{\prime},b^{\prime\prime}\}$ where $b^{\prime}$ is the neighbor of $p$ in $W\setminus P$ and $b^{\prime\prime}$ is the neighbor of $p^{\prime}$ in $W\setminus P$ (because $N_{H}(v)=\{a,b^{\prime}\}$ or $N_{H}(v)=\{a,b^{\prime\prime}\}$ would imply that $v$ and $H$ form a theta). Let $Q=p\hbox{-}\cdots\hbox{-}q$ be the sector of $W$ that contains $b^{\prime}$. By (3) applied to $v$ and $Q$, we have $ap\in E(G)$ and $b^{\prime\prime}q\in E(G)$. So, $b^{\prime\prime}$ is the unique neighbor of $v$ in the sector $R=p^{\prime}\hbox{-}\cdots\hbox{-}q$ of $W$. By (3) applied to $v$ and $R$, we have $ap^{\prime}\in E(G)$ and $b^{\prime}q\in E(G)$. So $H$ has length six, a contradiction.
We proved that for every sector $P$ of $W$, either $v$ has no neighbors in $P$, or $v$ has two neighbors in $P$, and those neighbors are adjacent. We may therefore assume that $v$ has neighbors in at least three distinct sectors of $W$, because if $v$ has neighbors in exactly two of them, then $H\cup\{v\}$ would be a prism. So, suppose that $P=p\hbox{-}\cdots\hbox{-}p^{\prime}$, $Q=q\hbox{-}\cdots\hbox{-}q^{\prime}$ and $R=r\hbox{-}\cdots\hbox{-}r^{\prime}$ are three distinct sectors of $W$, and $v$ is adjacent to $x,x^{\prime}\in P$, to $y,y^{\prime}\in Q$ and to $z,z^{\prime}\in R$. Suppose up to symmetry that $p$, $x$, $x^{\prime}$, $p^{\prime}$, $q$, $y$, $y^{\prime}$, $q^{\prime}$, $r$, $z$, $z^{\prime}$ and $r^{\prime}$ appear in this order along $H$. Then there is a theta in $G$ with ends $c,v$ and paths $v\hbox{-}x\hbox{-}P\hbox{-}p\hbox{-}c$, $v\hbox{-}y\hbox{-}Q\hbox{-}q\hbox{-}c$ and $v\hbox{-}z\hbox{-}R\hbox{-}r\hbox{-}c$, a contradiction. This proves (3).
To conclude the proof, suppose for a contradiction that the interiors of two distinct sectors of $W$ are contained in the same connected component of $G\setminus T$. Then there exists a path $Y=v\hbox{-}\cdots\hbox{-}w$ in $G\setminus T$ and two sectors $P=p\hbox{-}\cdots\hbox{-}p^{\prime}$ and $Q=q\hbox{-}\cdots\hbox{-}q^{\prime}$ of $W$ such that $v$ has neighbors in $P^{*}$ and $w$ has neighbors in $Q^{*}$. By (3), $v$ is anticomplete to $W\setminus P$ and $w$ is anticomplete to $W\setminus Q$ (in particular, $Y$ has length at least one). By choosing such a path $Y$ to be minimal, we deduce that $Y^{*}$ is anticomplete to $H$.
Suppose that $v$ has a unique neighbor, or two distinct and non-adjacent neighbors in $P$. Next, assume that $w$ has a neighbor $d$ in $H$ that is distinct from $b^{\prime}$ and $b^{\prime\prime}$ where $b^{\prime}$ is the neighbor of $p$ in $W\setminus P$ and $b^{\prime\prime}$ is the neighbor of $p^{\prime}$ in $W\setminus P$, then let $d^{\prime}$ be a neighbor of $c$ in $H\setminus(P\cup\{b^{\prime},b^{\prime\prime}\})$ ($d^{\prime}$ exists for otherwise, $W$ would be a prism or a special wheel). We choose $d$ and $d^{\prime}$ such that the path $R$ in $H\setminus P$ from $d$ to $d^{\prime}$ is minimal. We now see that if $v$ has a unique neighbor $a$ in $P$, then $P\cup Y\cup R\cup\{c\}$ contains a theta with ends $a$ and $c$, a contradiction. Also, if $v$ has two distinct non-adjacent neighbors in $P$, then $P\cup Y\cup R\cup\{c\}$ contains a theta with ends $c$ and $v$. So, $w$ has only two possible neighbors in $H$, namely, $b^{\prime}$ and $b^{\prime\prime}$. Due to symmetry, we may assume that $b^{\prime}w\in E(G)$ (so $b^{\prime\prime}w\notin E(G)$). It follows that $b^{\prime}$ is non-adjacent to $c$. If $v$ has a unique neighbor in $P$, then $H\cup Y$ is a theta in $G$, so $v$ has a neighbor in $P$ that is non-adjacent to $p$. In particular, there exists a path $R^{\prime}$ from $v$ to $p^{\prime}$ in $P\cup\{v\}$ that contains no neighbor of $p$. It follows that $R^{\prime}\cup Q\cup Y\cup\{c\}$ is a theta in $G$ with ends $b^{\prime}$ and $c$.
We deduce that $v$ has exactly two neighbors in $P$, and those neighbors are adjacent. By the same argument, we can prove that $w$ has exactly two neighbors in $P$ that are adjacent. But now $H\cup Y$ is a prism in $G$, a contradiction. This completes the proof of Theorem 3.1. ∎
Theorem 3.2.
Let $G$ be a (theta, prism)-free graph and let $W=(H,c)$ be a special wheel in $G$ whose long sectors have lengths at least three. Let $a^{\prime\prime},b^{\prime\prime}\in G\setminus N[Z(W)]$ belong to (the interiors of) distinct sectors of $W$. Then $N[Z(W)]$ separates $a^{\prime\prime}$ and $b^{\prime\prime}$ in $G$.
Proof.
Let $ab$ be the sector of length one of $W$ and let $d$ be the neighbor of $c$ in $H\setminus\{a,b\}$. Let $a^{\prime}$ be the neighbor of $d$ in the long sector of $W$ containing $a$ and let $b^{\prime}$ be the neighbors of $d$ in the long sector of $W$ containing $b$. Then $Z(W)=\{a,a^{\prime},b,b^{\prime},c,d\}$. Let $P$ be the path in $H\setminus d$ from $a$ to $a^{\prime}$ and let $Q$ be the path of $H\setminus d$ from $b$ to $b^{\prime}$. Assume, without loss of generality, that $a^{\prime\prime}\in P^{*}\setminus N[Z(W)]$ and let $b^{\prime\prime}\in Q^{*}\setminus N[Z(W)]$.
Let $T=N[c]\cup(N[\{a,b,a^{\prime},b^{\prime},d\}]\setminus H)$. Then $T\subseteq N[Z(W)]$, and so it suffices to show that $T$ separates $a^{\prime\prime}$ and $b^{\prime\prime}$ (note that $a^{\prime\prime},b^{\prime\prime}\notin T$). Suppose not. Then there exists a path $Y=v\hbox{-}\cdots\hbox{-}w$ in $G\setminus T$ such that $v$ has neighbors in $P^{*}$, $w$ has neighbors in $Q^{*}$, $Y\setminus v$ is anticomplete to $W\setminus P$ and $Y\setminus w$ is anticomplete to $W\setminus Q$ (note that possibly $v=w$).
Let $x$ be the neighbor of $v$ in $P$ closest to $a$ along $P$ and let $x^{\prime}$ be the neighbor of $v$ in $P$ closest to $a^{\prime}$ along $P$. Let $y$ be the neighbor of $w$ in $Q$ closest to $b$ along $Q$ and let $y^{\prime}$ be the neighbor of $w$ in $Q$ closest to $b^{\prime}$ along $Q$.
If $x=x^{\prime}$, then there is a theta in $G$ with ends $x$ and $d$ and paths $x\hbox{-}P\hbox{-}a^{\prime}\hbox{-}d$, $x\hbox{-}P\hbox{-}a\hbox{-}c\hbox{-}d$ and $x\hbox{-}v\hbox{-}Y\hbox{-}w\hbox{-}y^{\prime}\hbox{-}Q\hbox{-}b^{\prime}\hbox{-}d$. So, $x\neq x^{\prime}$, and symmetrically we have $y\neq y^{\prime}$. If $xx^{\prime}\notin E(G)$, then there is a theta in $G$ with ends $v$ and $d$ and paths $v\hbox{-}x^{\prime}\hbox{-}P\hbox{-}a^{\prime}\hbox{-}d$, $v\hbox{-}x\hbox{-}P\hbox{-}a\hbox{-}c\hbox{-}d$ and $v\hbox{-}Y\hbox{-}w\hbox{-}y^{\prime}\hbox{-}Q\hbox{-}b^{\prime}\hbox{-}d$. So, $xx^{\prime}\in E(G)$, and symmetrically we can prove that $yy^{\prime}\in E(G)$. But now $H\cup Y$ is a prism in $G$, a contradiction. This completes the proof of Theorem 3.2. ∎
4. Breaking a pyramid
A pyramid is a graph $\Sigma$ consisting of a vertex $a$, a triangle $\{b_{1},b_{2},b_{3}\}$ disjoint from $a$ and three paths $P_{1},P_{2},P_{3}$ in $\Sigma$ of length at least two, such that for each $i\in[3]$, the ends of $P_{i}$ are $a$ and $b_{i}$, and for all distinct $i,j\in[3]$, the sets $V(P_{i})\setminus\{a\}$ and $V(P_{j})\setminus\{a\}$ are disjoint, $b_{i}b_{j}$ is the only edge of $G$ with an end in $V(P_{i})\setminus\{a\}$ and an end in $V(P_{j})\setminus\{a\}$, and for every $i\neq j\in\{1,2,3\}$ $P_{i}\cup P_{j}$ is a hole (the assumption that $P_{1},P_{2},P_{3}$ have length at least two is non-standard; usually, one of the paths is allowed to have length $1$, and our definition above would refer to a “long” pyramid.)
We say that $a$ is the apex of $\Sigma$, the triangle $\{b_{1},b_{2},b_{3}\}$ is the base of $\Sigma$, and $P_{1},P_{2},P_{3}$ are the paths of $\Sigma$. We also define $Z(\Sigma)=N_{\Sigma}[a]\cup\{b_{1},b_{2},b_{3}\}$ (so we have $|Z(\Sigma)|=7$). For a graph $G$, by a pyramid in $G$ we mean an induced subgraph of $G$ which is a pyramid (see Figure 3).
Figure 3. A pyramid $\Sigma$. Dashed lines represent paths of arbitrary (possibly zero) length, and circled nodes represent the vertices in $Z(\Sigma)$.
The main result of this section, Theorem 4.1 below, follows from much more general results of [tw8]. However, there is also a short and self-contained proof, which we include here:
Theorem 4.1.
Let $G$ be a (theta, prism)-free graph and let $\Sigma$ be a pyramid in $G$ with apex $a$, base $\{b_{1},b_{2},b_{3}\}$ and paths $P_{1},P_{2}$ and $P_{3}$ as in the definition. Let $u,v\in G\setminus N[Z(\Sigma)]$ belong to distinct paths of $\Sigma$. Then $N[Z(\Sigma)]$ separates $u$ and $v$ in $G$.
Proof.
Suppose not. Then there exist $u,v\in G\setminus N[Z(\Sigma)]$, belonging to distinct paths of $\Sigma$, such that $N[Z(\Sigma)]$ does not separate $u$ and $v$ in $G$. It follows that for distinct $i,j\in[3]$, there exists a path $Q=x\hbox{-}\cdots\hbox{-}y$ in $G\setminus(\Sigma\cup N[Z(\Sigma)])$ such that $x$ has a neighbor in $P^{*}_{i}$ and $y$ has a neighbor in $P^{*}_{j}$. We choose $i,j\in[3]$ and $Q$ subject to the minimality of $Q$. By symmetry, we may assume that $i=1$ and $j=2$.
From the minimality of $Q$ and the fact that $Q\subseteq V(G)\setminus(\Sigma\cup N[Z(\Sigma)])$, it follows that:
  • $N_{P_{1}}(x)\subseteq P_{1}\setminus Z(\Sigma)$, and $Q\setminus x$ and $P_{1}$ are anticomplete in $G$.
  • $N_{P_{2}}(y)\subseteq P_{2}\setminus Z(\Sigma)$, and $Q\setminus y$ and $P_{2}$ are anticomplete in $G$.
Now, if some vertex of $Q$ has a neighbor in $P_{3}$, then by the minimality of $Q$, we must have $x=y$. In particular, $x$ has neighbors in $P_{1}$, $P_{2}$ and $P_{3}$. Since $a$ and $x$ are not adjacent in $G$ (for otherwise there is a theta in $G$), it follows that the three paths in $G$ from $a$ to $x$ with interiors in $P_{1}$, $P_{2}$ and $P_{3}$ form a theta in $G$ with ends $a$ and $x$, a contradiction. We deduce that $Q$ and $P_{3}$ are anticomplete in $G$.
Let $x^{\prime}$ be the neighbor of $x$ in $P_{1}$ closest to $a$ along $P_{1}$ and let $x^{\prime\prime}$ be the neighbor of $x$ in $P_{1}$ closest to $b_{1}$ along $P_{1}$. Similarly, let $y^{\prime}$ be the neighbor of $y$ in $P_{2}$ closest to $a$ along $P_{2}$ and let $y^{\prime\prime}$ be the neighbor of $y$ in $P_{2}$ closest to $b_{2}$ along $P_{2}$. Recall that $x^{\prime},x^{\prime\prime}\in P_{1}\setminus Z(\Sigma)$ and $y^{\prime},y^{\prime\prime}\in P_{2}\setminus Z(\Sigma)$. If $x^{\prime}=x^{\prime\prime}$, then there is a theta in $G$ with ends $a,x^{\prime}$ and paths $a\hbox{-}P_{1}\hbox{-}x^{\prime}$, $a\hbox{-}P_{2}\hbox{-}y^{\prime}\hbox{-}y\hbox{-}Q\hbox{-}x\hbox{-}x^{\prime}$ and $a\hbox{-}P_{3}\hbox{-}b_{3}\hbox{-}b_{1}\hbox{-}P_{1}\hbox{-}x^{\prime}$. Also, if $x^{\prime}$ and $x^{\prime\prime}$ are distinct and adjacent in $G$, then there is a prism in $G$ with triangles $x^{\prime\prime}xx^{\prime}$ and $b_{1}b_{2}b_{3}$ and paths $x^{\prime\prime}\hbox{-}P_{1}\hbox{-}b_{1}$, $x\hbox{-}Q\hbox{-}y\hbox{-}y^{\prime\prime}\hbox{-}P_{2}\hbox{-}b_{2}$ and $x^{\prime}\hbox{-}P_{1}\hbox{-}a\hbox{-}P_{3}\hbox{-}b_{3}$. Hence, we have $x^{\prime}\neq x^{\prime\prime}$ and $x^{\prime}x^{\prime\prime}\notin E(G)$. But now there is a theta in $G$ with ends $a,x$ and paths $a\hbox{-}P_{1}\hbox{-}x^{\prime}\hbox{-}x$, $a\hbox{-}P_{2}\hbox{-}y^{\prime}\hbox{-}y\hbox{-}Q\hbox{-}x$ and $a\hbox{-}P_{3}\hbox{-}b_{3}\hbox{-}b_{1}\hbox{-}P_{1}\hbox{-}x^{\prime\prime}\hbox{-}x$, a contradiction. This completes the proof of Theorem 4.1. ∎
5. Alignments and Connectifiers
This section covers a number of definitions and a result from [tw15], which we will use in the proof of Theorem 2.1.
Let $G$ be a graph, let $P$ be a path in $G$ and let $X\subseteq V(G)\setminus P$. We say that $(P,X)$ is an alignment if every vertex of $X$ has at least one neighbor in $P$ and one may write $P=p_{1}\hbox{-}\cdots\hbox{-}p_{n}$ and $X=\{x_{1},\ldots,x_{k}\}$ for $k,n\in\mathbb{N}$ such that there exist $1\leq i_{1}\leq j_{1}<i_{2}\leq j_{2}<\cdots<i_{k}\leq j_{k}\leq n$ where $N_{P}(x_{l})\subseteq p_{i_{l}}\hbox{-}P\hbox{-}p_{j_{l}}$ for every $l\in[k]$. This is a little different from the definition in [tw15], but the difference is not substantial, and using this definition is more convenient for us here. In this case, we say that $x_{1},\ldots,x_{k}$ is the order on $X$ given by the alignment $(P,X)$. An alignment $(P,X)$ is wide if each of $x_{1},\ldots,x_{k}$ has two non-adjacent neighbors in $P$, spiky if each of $x_{1},\ldots,x_{k}$ has a unique neighbor in $P$ and triangular if each of $x_{1},\ldots,x_{k}$ has exactly two neighbors in $P$ and those neighbors are adjacent. An alignment is consistent if it is wide, spiky or triangular. See Figure 4.
Figure 4. A consistent alignment which is spiky (top left), triangular (top right) and wide (bottom).
By a caterpillar we mean a tree $C$ with maximum degree three such that no two branch vertices in $C$ are adjacent, and such that there exists a path $P$ in $C$ containing all branch vertices of $C$. We call a minimal such path $P$ the spine of $C$. (We note that our definition of a “caterpillar” is non-standard in multiple ways.) By a subdivided star we mean a graph isomorphic to a subdivision of the complete bipartite graph $K_{1,\delta}$ for some $\delta\geq 3$. In other words, a subdivided star is a tree with exactly one branch vertex, which we call its root. For a graph $H$, a vertex $v$ of $H$ is said to be simplicial if $N_{H}(v)$ is a clique. We denote by $\mathcal{Z}(H)$ the set of all simplicial vertices of $H$. Note that for every tree $T$, $\mathcal{Z}(T)$ is the set of all leaves of $T$. An edge $e$ of a tree $T$ is said to be a leaf-edge of $T$ if $e$ is incident with a leaf of $T$. It follows that if $H$ is the line graph of a tree $T$, then $\mathcal{Z}(H)$ is the set of all vertices in $H$ corresponding to the leaf-edges of $T$.
Let $H$ be a graph that is either a caterpillar, or the line graph of a caterpillar, or a subdivided star with root $r$, or the line graph of a subdivided star with root $r$. We define an induced subgraph of $H$, denoted by $P(H)$, which we will use throughout the paper. If $H$ is a path (possibly of length zero), then let $P(H)=H$. If $H$ is a caterpillar, then let $P(H)$ be the spine of $H$. If $H$ is the line graph of a caterpillar $C$, then let $P(H)$ be the path in $H$ consisting of the vertices of $H$ that correspond to the edges of the spine of $C$. If $H$ is a subdivided star with root $r$, then let $P(H)=\{r\}$. If $H$ is the line graph of a subdivided star $S$ with root $r$, let $P(H)$ be the clique of $H$ consisting of the vertices of $H$ that correspond to the edges of $S$ incident with $r$. The legs of $H$ are the components of $H\setminus P(H)$. Let $G$ be a graph and let $H$ be an induced subgraph of $G$ that is either a caterpillar, or the line graph of a caterpillar, or a subdivided star or the line graph of a subdivided star. Let $X\subseteq V(G)\setminus H$ such that every vertex of $X$ has a unique neighbor in $H$ and $N_{H}(X)=\mathcal{Z}(H)$ (see Figure 5). We call $(H,X)$ a connectifier. Also, if $H$ is a single vertex and $X\subseteq N(H)$, we call $(H,X)$ a connectifier as well. We say that the connectifier $(H,X)$ is concentrated if $H$ is a subdivided star or the line graph of a subdivided star or a singleton.
Let $(H,X)$ be a connectifier in $G$ which is not concentrated. So $H$ is a caterpillar or the line graph of a caterpillar. Let $S$ be the set of vertices of $H\setminus P(H)$ that have neighbors in $P(H)$. Then $(P(H),S)$ is an alignment. Let $s_{1},\ldots,s_{k}$ be the corresponding order on $S$ given by $(P(H),S)$. Now, order the vertices of $X$ as $x_{1},\ldots,x_{k}$ where for every $i\in[k]$, the vertex $x_{i}$ has a neighbor in the leg of $H$ containing $s_{i}$. We say that $x_{1},\ldots,x_{k}$ is the order on $X$ given by $(H,X)$.
The following was proved in [tw15]:
Theorem 5.1 (Chudnovsky, Gartland, Hajebi, Lokshtanov and Spirkl; Theorem 5.2 in [tw15]).
[…]
  • $(H,S^{\prime})$ is a connectifier, or
  • $H$ is a path and every vertex in $S^{\prime}$ has a neighbor in $H$.
Figure 5. Examples of a connectifier. Circled nodes represent the vertices in $X$.
6. Amiability
The two notions of “amiability” and “amicability,” first introduced in [tw15], are at the heart of the proof of Theorem 2.1. We deal with the former in this section and leave the latter for the next one.
Let $s\in\mathbb{N}$ and let $G$ be a graph. An $s$-trisection in $G$ is a separation $(D_{1},Y,D_{2})$ in $G$ such that the following hold.
  • $Y$ is a stable set with $|Y|=s$.
  • $D_{1}$ and $D_{2}$ are components of $G\setminus Y$ with $N(D_{1})=N(D_{2})=Y$.
  • $D_{1}$ is a path and for every $y\in Y$ there exists $d_{y}\in D_{1}$ such that $N_{Y}(d_{y})=\{y\}$.
(The reader may notice that we will never use the second condition in the third bullet point. It was however necessary in [tw15], so we keep it for easier cross-referencing.)
We say that a graph class $\mathcal{G}$ is amiable if there is a function $\sigma:\mathbb{N}\rightarrow\mathbb{N}$ with the following property. Let $x\in\mathbb{N}$, let $G\in\mathcal{G}$ and let $(D_{1},Y,D_{2})$ be a $\sigma(x)$-trisection in $G$. Then there exist $H\subseteq D_{2}$ and $X\subseteq Y$ with $|X|=x$ such that the following hold.
  • $(D_{1},X)$ is a consistent alignment.
  • $(H,X)$ is either a connectifier or a consistent alignment.
  • If $(H,X)$ is not a concentrated connectifier, then the orders given on $X$ by $(D_{1},X)$ and by $(H,X)$ are the same.
In this case, we say that $H$ and $X$ are given by amiability. The main result of this section is the following:
Theorem 6.1.
For every $t\in\mathbb{N}$, the class $\mathcal{C}_{t}$ is amiable. Moreover, with notation as in the definition of amiability, if $(H,X)$ is a connectifier, then we have $|H|>1$.
In order to prove Theorem 6.1, first we prove the following lemma:
Lemma 6.2.
Let $d,s\in\mathbb{N}$, let $G$ be a theta-free graph and let $Y$ be a stable set in $G$ of cardinality $3s(d+1)$. Let $P$ be a path in $G\setminus Y$ such that every vertex in $Y$ has a neighbor in $P$, and each vertex of $P$ has fewer than $d$ neighbors in $Y$. Assume that for every two vertices $y,y^{\prime}\in Y$, there is a path $R$ in $G$ from $y$ to $y^{\prime}$ such that $P$ and $R^{*}$ are disjoint and anticomplete in $G$. Then there is an $s$-subset $S$ of $Y$ such that $(P,S)$ is a consistent alignment.
Proof.
For every vertex $y\in Y$, let $P_{y}$ be the (unique) path in $P$ with the property that $y$ is complete to the ends of $P_{y}$ and anticomplete to $P\setminus P_{y}$. Let $I$ be the graph with $V(I)=Y$ such that two distinct vertices $y,y^{\prime}\in Y$ are adjacent in $I$ if and only if $P_{y}\cap P_{y^{\prime}}\neq\emptyset$. Then $I$ is an interval graph and so $I$ is perfect [Golumbic]. Since $|V(I)|=3s(d+1)$, it follows that $I$ contains either a clique of cardinality $d+1$ or a stable set of cardinality $3s$.
Assume that $I$ contains a clique of cardinality $d+1$. Then there exists $C\subseteq Y$ with $|C|=d+1$ and $p\in P$ such that $p\in P_{y}$ for every $y\in C$. Since $p\in P$ has fewer than $d$ neighbors in $C\subseteq Y$, it follows that there are at least two vertices $y,y^{\prime}\in C\setminus N(p)$. Since $p\in P_{y}\cap P_{y^{\prime}}$, it follows that $P\setminus\{p\}$ has two components, and each of $y$ and $y^{\prime}$ has a neighbor in each component of $P\setminus\{p\}$. It follows that there are two paths $P_{1}$ and $P_{2}$ from $y$ to $y^{\prime}$ with disjoint and anticomplete interiors contained in $P$. On the other hand, there is a path $R$ in $G$ from $y$ to $y^{\prime}$ such that $P$ and $R^{*}$ are disjoint and anticomplete in $G$. It follows that $P_{1},P_{2}$ and $R$ are pairwise internally disjoint and anticomplete. But now there is a theta in $G$ with ends $y,y^{\prime}$ and paths $P_{1},P_{2},R$, a contradiction.
We deduce that $I$ contains a stable set $S^{\prime}$ of cardinality $3s$. From the definition of $I$, it follows that $(P,S^{\prime})$ is an alignment. Hence, since every vertex in $S^{\prime}$ has one, two adjacent, or at least two non-adjacent neighbors in $P$, there exists $S\subseteq S^{\prime}\subseteq Y$ with $|S|=s$ such that $(P,S)$ is a consistent alignment. This completes the proof of Lemma 6.2. ∎
Proof of Theorem 6.1.
For every $x\in\mathbb{N}$, let
[…]
and let
$\sigma(x)=3s(t+1).$
We will show that $\mathcal{C}_{t}$ is amiable with respect to $\sigma:\mathbb{N}\rightarrow\mathbb{N}$ as defined above. Let $x\in\mathbb{N}$, let $G\in\mathcal{C}_{t}$ and let $(D_{1},Y,D_{2})$ be a $\sigma(x)$-trisection in $G$. Then $Y$ is a stable set of cardinality $3s(t+1)$, $D_{1}$ is a path in $G\setminus Y$ and every vertex in $Y$ has a neighbor in $D_{1}$. Moreover, since $G$ is $K_{1,t}$-free, no vertex in $D_{1}$ has $t$ or more neighbors in $Y$, and since $N(D_{2})=Y$, it follows that for every two vertices $y,y^{\prime}\in Y$, there is a path $R$ in $G$ from $y$ to $y^{\prime}$ with $R^{*}\subseteq D_{2}$, and so $D_{1}$ and $R^{*}$ are disjoint and anticomplete in $G$. By Lemma 6.2, there exists $S\subseteq Y$ with $|S|=s$ such that $(D_{1},S)$ is a consistent alignment.
[…]
  • $(H_{2},S^{\prime})$ is a connectifier.
  • $H_{2}$ is a path and every vertex of $S^{\prime}$ has a neighbor in $H_{2}$.
First, assume that $(H_{2},S^{\prime})$ is a concentrated connectifier. Then, since $|S^{\prime}|\geq t$ and $G$ is $K_{1,t}$-free, it follows that $|H_{2}|>1$. Now, since $|S^{\prime}|\geq x$, we may choose a concentrated connectifier $(H,X)$ where $X$ is an $x$-subset of $S^{\prime}\subseteq S\subseteq Y$ and $H$ is an induced subgraph $H_{2}\subseteq D_{2}$ with $|H|>1$. In particular, $H$ and $X$ satisfy the definition of amiability.
Next, assume that $(H_{2},S^{\prime})$ is a connectifier which is not concentrated. Consider the orders on $S^{\prime}$ given by $(D_{1},S^{\prime})$ and by $(H_{2},S^{\prime})$. Since $|S^{\prime}|\geq x^{2}$, it follows from the Erdős-Szekeres theorem [ES] that there is an $x$-subset $X$ of $S^{\prime}\subseteq S\subseteq Y$ as well as an induced subgraph $H$ of $H_{2}\subseteq D_{2}$ such that:
  • $(D_{1},X)$ is a consistent alignment (because $(D_{1},S)$ is);
  • $(H,X)$ is a connectifier which is not concentrated; and
  • The orders given on $X$ by $(D_{1},X)$ and by $(H,X)$ are the same.
It follows that $H$ and $X$ satisfy the definition of amiability.
Finally, assume that $H_{2}$ is a path and every vertex in $S^{\prime}$ has a neighbor in $H_{2}$. Let $H=H_{2}$. Recall that $(D_{1},S^{\prime})$ is an alignment. In particular, $S^{\prime}$ is a stable set of cardinality $3x^{2}(t+1)$, and since $G$ is $K_{1,t}$-free, no vertex in $H_{2}$ has $t$ or more neighbors in $S^{\prime}$. Also, for every two vertices $y,y^{\prime}\in S$, there is a path $R$ in $G$ from $y$ to $y^{\prime}$ such that $R^{*}\subseteq D_{1}$, and so $H$ and $R^{*}$ are disjoint and anticomplete in $G$. By Lemma 6.2, there exists $S^{\prime\prime}\subseteq S^{\prime}\subseteq S$ with $|S^{\prime\prime}|=x^{2}$ such that $(H,S^{\prime\prime})$ is a consistent alignment. Consider the order on $S^{\prime\prime}$ given by $(D_{1},S^{\prime\prime})$ and by $(H,S^{\prime\prime})$. Since $|S^{\prime\prime}|=x^{2}$, it follows from the Erdős-Szekeres theorem [ES] that there is an $x$-subset $X$ of $S^{\prime\prime}\subseteq S^{\prime}\subseteq S\subseteq Y$ such that such that:
  • $(D_{1},X)$ is a consistent alignment (because $(D_{1},S)$ is);
  • $(H,X)$ is a consistent alignment (because $(H,S^{\prime\prime})$ is); and
  • The orders given on $X$ by $(D_{1},X)$ and by $(H,X)$ are the same.
So $H$ and $X$ satisfy the definition of amiability. This completes the proof of Theorem 6.1
7. Amicability
Here we complete the proof of Theorem 2.1, beginning with the following definition.
Let $m\in\mathbb{N}$ and let $\mathcal{G}$ be a graph class. We say that $\mathcal{G}$ is $m$-amicable if $\mathcal{G}$ is amiable and the following holds. Let $\sigma:\mathbb{N}\rightarrow\mathbb{N}$ be as in the definition of amiability for $\mathcal{G}$. Let $G\in\mathcal{G}$ and let $(D_{1},Y,D_{2})$ be a $\sigma(7)$-trisection in $G$. Let $X=\{x_{1},\ldots,x_{7}\}\subseteq Y$ be given by amiability such that $x_{1},\ldots,x_{7}$ is the order on $X$ given by $(D_{1},X)$. Let $D_{1}=d_{1}\hbox{-}\cdots\hbox{-}d_{k}$ such that traversing $D_{1}$ from $d_{1}$ to $d_{k}$, the first vertex in $D_{1}$ with a neighbor in $X$ is a neighbor of $x_{1}$. Let $i\in[k]$ be maximum such that $x_{1}$ is adjacent to $d_{i}$ and let $j\in[k]$ be minimum such that $x_{7}$ is adjacent to $d_{j}$. Then there exists a subset $Z$ of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$ with $|Z|\leq m$ such that $N[Z]$ separates $d_{i}$ and $d_{j}$. It follows that $N[Z]$ separates $d_{1}\hbox{-}D_{1}\hbox{-}d_{i}$ and $d_{j}\hbox{-}D_{1}\hbox{-}d_{k}$ (see Figure 6).
Figure 6. Amicability – Note that $Z$ is contained in the highlighted set.
Figure 7. $H$ is a caterpillar. Circled nodes depict the vertices in $Z(\Sigma)$.
We prove that:
Theorem 7.1.
For every $t\in\mathbb{N}$, the class $\mathcal{C}_{t}$ is $\max\{2t,7\}$-amicable.
Proof.
By Theorem 6.1, $\mathcal{C}_{t}$ is amiable, and with notation as in the definition of amiability, if $(H,X)$ is a connectifier, then we have $|H|>1$. Let $\sigma:\mathbb{N}\rightarrow\mathbb{N}$ be as in the definition of amiability for $\mathcal{C}_{t}$. Let $G\in\mathcal{C}_{t}$ and let $(D_{1},Y,D_{2})$ be a $\sigma(7)$-trisection in $G$. Let $X=\{x_{1},\ldots,x_{7}\}\subseteq Y$ be given by amiability such that $x_{1},\ldots,x_{7}$ is the order on $X$ given by the consistent alignment $(D_{1},X)$. Let $D_{1}=d_{1}\hbox{-}\cdots\hbox{-}d_{k}$ and $i,j\in[k]$ be as in the definition of amicability. Our goal is to show that there exists a subset $Z$ of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$ with $|Z|\leq\max\{2t,7\}$ such that $N[Z]$ separates $d_{i}$ and $d_{j}$.
Let $i^{\prime}\in[k]$ be minimum such that $x_{4}$ is adjacent to $d_{i^{\prime}}$, let $j^{\prime}\in[k]$ be maximum such that $x_{4}$ is adjacent to $d_{j^{\prime}}$, and let $H$ be the induced subgraph of $D_{2}$ given by amiability. It follows that $i+2<i^{\prime}\leq j^{\prime}<j-2$, $(H,X)$ is either a connectifier with $|H|>1$ or a consistent alignment, and if $(H,X)$ is not a concentrated connectifier, then $x_{1},\ldots,x_{7}$ is the order on $X$ given by $(H,X)$. When $(H,X)$ is a connectifier with $|H|>1$, then for each $l\in[7]$, let $r_{l}$ be the unique neighbor of $x_{l}$ in $H$ (so $r_{l}\in\mathcal{Z}(H)$) and let $L_{l}$ be the (unique) shortest path in $H$ from $r_{l}$ to a vertex $s_{l}\in N_{H}[P(H)]$. It follows that $s_{l}\in H\setminus P(H)$ unless $H$ is the line graph of a subdivided star where not all edges of the star are subdivided, in which case we have $r_{l}=s_{l}\in P(H)=\mathcal{Z}(H)=H$.
First, consider the case where $H$ is a caterpillar. It follows that for each $l\in[7]$, we have $s_{l}\in H\setminus P(H)$ and $s_{l}$ has a unique neighbor $p_{l}\in P(H)$. Since $G$ is theta-free, it follows that $(D_{1},X)$ is triangular, and so $j^{\prime}=i^{\prime}+1$ (see Figure 7). Let $\Sigma$ be the pyramid with apex $p_{4}$, base $\{d_{i^{\prime}},x_{4},d_{j^{\prime}}\}$ and paths
$P_{1}=p_{4}\hbox{-}P(H)\hbox{-}p_{1}\hbox{-}s_{1}\hbox{-}L_{1}\hbox{-}r_{1}\hbox{-}x_{1}\hbox{-}d_{i}\hbox{-}D_{1}\hbox{-}d_{i^{\prime}};$
$P_{2}=p_{4}\hbox{-}s_{4}\hbox{-}L_{4}\hbox{-}r_{4}\hbox{-}x_{4};$
$P_{3}=p_{4}\hbox{-}P(H)\hbox{-}p_{7}\hbox{-}s_{7}\hbox{-}L_{7}\hbox{-}r_{7}\hbox{-}x_{7}\hbox{-}d_{j}\hbox{-}D_{1}\hbox{-}d_{j^{\prime}}.$
Then $Z(\Sigma)$ is a $7$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$. Moreover, we have $d_{i}\in P_{1}^{*}\setminus N[Z(\Sigma)]$ and $d_{j}\in P_{3}^{*}\setminus N[Z(\Sigma)]$. Therefore, by Theorem 4.1, $N[Z(\Sigma)]$ separates $d_{i}$ and $d_{j}$, as desired.
Figure 8. $H$ is the line graph of a caterpillar and $(D_{1},X)$ is spiky. Circled nodes represent the vertices in $Z(\Sigma)$.
Figure 9. $H$ is the line graph of a caterpillar and $(D_{1},X)$ is wide. Circled nodes represent the vertices in $Z(\Sigma)$.
Second, consider the case where $H$ is the line graph of a caterpillar. It follows that for each $l\in[7]$, we have $s_{l}\in H\setminus P(H)$ and $s_{l}$ has exactly two neighbors $p_{l},q_{l}\in P(H)$, where $p_{l}$ and $q_{l}$ are adjacent, and the vertices $p_{1},q_{1},p_{2},q_{2},\ldots,p_{7},q_{7}$ appear on $P(H)$ in this order. Since $G$ is prism-free, it follows that $(D_{1},X)$ is either spiky or wide. Suppose that $(D_{1},X)$ is spiky (see Figure 8). Then $i^{\prime}=j^{\prime}$. Let $\Sigma$ be the pyramid with apex $d_{i^{\prime}}=d_{j^{\prime}}$, base $\{p_{4},s_{4},q_{4}\}$ and paths
$P_{1}=d_{i^{\prime}}\hbox{-}D_{1}\hbox{-}d_{i}\hbox{-}x_{1}\hbox{-}r_{1}\hbox{-}L_{1}\hbox{-}s_{1}\hbox{-}q_{1}\hbox{-}P(H)\hbox{-}p_{4};$
$P_{2}=d_{i^{\prime}}\hbox{-}x_{4}\hbox{-}r_{4}\hbox{-}L_{4}\hbox{-}s_{4};$
$P_{3}=d_{i^{\prime}}\hbox{-}D_{1}\hbox{-}d_{j}\hbox{-}x_{7}\hbox{-}r_{7}\hbox{-}L_{7}\hbox{-}s_{7}\hbox{-}p_{7}\hbox{-}P(H)\hbox{-}q_{4}.$
Then $Z(\Sigma)$ is a $7$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$. Moreover, we have $d_{i}\in P_{1}^{*}\setminus N[Z(\Sigma)]$ and $d_{j}\in P_{3}^{*}\setminus N[Z(\Sigma)]$. So by Theorem 4.1, $N[Z(\Sigma)]$ separates $d_{i}$ and $d_{j}$. Now assume that $(D_{1},X)$ is wide (see Figure 9). Then $j^{\prime}-i^{\prime}>1$. Let $\Sigma$ be the pyramid with apex $x_{4}$, base $\{p_{4},s_{4},q_{4}\}$ and paths
$P_{1}=x_{4}\hbox{-}d_{i^{\prime}}\hbox{-}D_{1}\hbox{-}d_{i}\hbox{-}x_{1}\hbox{-}r_{1}\hbox{-}L_{1}\hbox{-}s_{1}\hbox{-}q_{1}\hbox{-}P(H)\hbox{-}p_{4};$
$P_{2}=x_{4}\hbox{-}r_{4}\hbox{-}L_{4}\hbox{-}s_{4};$
$P_{3}=x_{4}\hbox{-}d_{j^{\prime}}\hbox{-}D_{1}\hbox{-}d_{j}\hbox{-}x_{7}\hbox{-}r_{7}\hbox{-}L_{7}\hbox{-}s_{7}\hbox{-}p_{7}\hbox{-}P(H)\hbox{-}q_{4}.$
Let $Z=(N(x_{4})\cap\Sigma)\cup\{p_{4},s_{4},q_{4}\}$. Then $Z(\Sigma)$ is a $7$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$. Also, we have $d_{i}\in P_{1}^{*}\setminus N[Z(\Sigma)]$ and $d_{j}\in P_{3}^{*}\setminus N[Z(\Sigma)]$. So by Theorem 4.1, $N[Z(\Sigma)]$ separates $d_{i}$ and $d_{j}$, as required.
Figure 10. $H$ is a subdivided star. Circled nodes represent the vertices in $Z(\Sigma)$.
Figure 11. $H$ is the line graph of a subdivided star and $(D_{1},X)$ is spiky. Circled nodes represent the vertices in $Z(\Sigma)$, and the highlighted path may be of length zero.
Third, consider the case where $H$ is a subdivided star with root $r$. It follows that $P(H)=\{r\}$ and $H\neq\{r\}$ (because $|H|>1$). For each $l\in[7]$, we have $r_{l},s_{l}\in H\setminus P(H)$ and $r_{l}$ is a leaf of $H$. Since $G$ is theta-free, it follows that $(D_{1},X)$ is triangular and so $j^{\prime}-i^{\prime}=1$ (see Figure 10). Let $\Sigma$ be the pyramid with apex $r$, base $\{d_{i^{\prime}},x_{4},d_{j^{\prime}}\}$ and paths
$P_{1}=r\hbox{-}s_{1}\hbox{-}L_{1}\hbox{-}r_{1}\hbox{-}x_{1}\hbox{-}d_{i}\hbox{-}D_{1}\hbox{-}d_{i^{\prime}};$
$P_{2}=r\hbox{-}s_{4}\hbox{-}L_{4}\hbox{-}r_{4}\hbox{-}x_{4};$
$P_{3}=r\hbox{-}s_{7}\hbox{-}L_{7}\hbox{-}r_{7}\hbox{-}x_{7}\hbox{-}d_{j}\hbox{-}D_{1}\hbox{-}d_{j^{\prime}}.$
Then $Z(\Sigma)$ is a $7$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$. Also, we have $d_{i}\in P_{1}^{*}\setminus N[Z(\Sigma)]$ and $d_{j}\in P_{3}^{*}\setminus N[Z(\Sigma)]$. So it follows from Theorem 4.1 that $N[Z(\Sigma)]$ separates $d_{i}$ and $d_{j}$, as desired.
Figure 12. $H$ is the line graph of a subdivided star, $(D_{1},X)$ is wide and the vertices $r_{4},s_{4},h_{4}$ are not all the same. Circled nodes represent the vertices in $Z(\Sigma)$, and the highlighted path has length at least one.
Figure 13. $H$ is the line graph of a subdivided star, $(D_{1},X)$ is wide and $r_{4}=s_{4}=h_{4}$. The hole $C$ is highlighted, and circled nodes represent the vertices in $Z(W)$.
Fourth, consider the case where $H$ is the line graph of a subdivided star. It follows that for each $l\in[7]$, either we have $s_{l}\in P(H)$, in which case we set $h_{l}=s_{l}$, or we have $s_{l}\in H\setminus P(H)$, in which case we choose $h_{l}$ to be the unique neighbor of $s_{l}$ in $P(H)$. Since $G$ is prism-free, it follows that $(D_{1},X)$ is either spiky or wide. There are now three cases to analyze:
Figure 14. One of $(D_{1},X)$ and $(H,X)$ is spiky and the other is triangular. The hole $C$ is highlighted, and circled nodes represent the vertices in $Z(W)$.
Figure 15. One of $(D_{1},X)$ and $(H,X)$ is wide. The hole $C$ is highlighted, and circled nodes represent the vertices in $Z(W)$.
  • Case 1.
    Suppose that $(D_{1},X)$ is spiky (see Figure 11). Then we have $i^{\prime}=j^{\prime}$. Consider the pyramid $\Sigma$ in $G$ with apex $d_{i^{\prime}}=d_{j^{\prime}}$, base $\{h_{1},h_{4},h_{7}\}$ and paths
    $P_{1}=d_{i^{\prime}}\hbox{-}D_{1}\hbox{-}d_{i}\hbox{-}x_{1}\hbox{-}r_{1}\hbox{-}L_{1}\hbox{-}s_{1}\hbox{-}h_{1};$
    $P_{2}=d_{i^{\prime}}\hbox{-}x_{4}\hbox{-}r_{4}\hbox{-}L_{4}\hbox{-}s_{4}\hbox{-}h_{4};$
    $P_{3}=d_{i^{\prime}}\hbox{-}D_{1}\hbox{-}d_{j}\hbox{-}x_{7}\hbox{-}r_{7}\hbox{-}L_{7}\hbox{-}s_{7}\hbox{-}h_{7}.$
    Then $Z(\Sigma)$ is a $7$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$. Moreover, we have $d_{i}\in P_{1}^{*}\setminus N[Z(\Sigma)]$ and $d_{j}\in P_{3}^{*}\setminus N[Z(\Sigma)]$. Thus, by Theorem 4.1, $N[Z(\Sigma)]$ separates $d_{i}$ and $d_{j}$.
  • Case 2.
    Suppose that $(D_{1},X)$ is wide and the vertices $r_{4},s_{4},h_{4}$ are not all the same (see Figure 12). Then $j^{\prime}-i^{\prime}>1$. Let $\Sigma$ be the pyramid with apex $x_{4}$, base $\{h_{1},h_{4},h_{7}\}$ and paths
    $P_{1}=x_{4}\hbox{-}d_{i^{\prime}}\hbox{-}D_{1}\hbox{-}d_{i}\hbox{-}x_{1}\hbox{-}r_{1}\hbox{-}L_{1}\hbox{-}s_{1}\hbox{-}h_{1};$
    $P_{2}=x_{4}\hbox{-}r_{4}\hbox{-}L_{4}\hbox{-}s_{4}\hbox{-}h_{4};$
    $P_{3}=x_{4}\hbox{-}d_{j^{\prime}}\hbox{-}D_{1}\hbox{-}d_{j}\hbox{-}x_{7}\hbox{-}r_{7}\hbox{-}L_{7}\hbox{-}s_{7}\hbox{-}h_{7}.$
    Then $Z(\Sigma)$ is a $7$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$, and we have $d_{i}\in P_{1}^{*}\setminus N[Z(\Sigma)]$ and $d_{j}\in P_{3}^{*}\setminus N[Z(\Sigma)]$. It follows from Theorem 4.1 that $N[Z(\Sigma)]$ separates $d_{i}$ and $d_{j}$.
  • Case 3.
    Suppose that $(D_{1},X)$ is wide and $r_{4}=s_{4}=h_{4}$ (see Figure 13). Then $j^{\prime}-i^{\prime}>1$. Let $C=x_{4}\hbox{-}d_{i^{\prime}}\hbox{-}D_{1}\hbox{-}d_{i}\hbox{-}x_{1}\hbox{-}r_{1}\hbox{-}L_{1}\hbox{-}s_{1}\hbox{-}h_{1}\hbox{-}h_{7}\hbox{-}s_{7}\hbox{-}L_{7}\hbox{-}r_{7}\hbox{-}x_{7}\hbox{-}d_{j}\hbox{-}D_{1}\hbox{-}d_{j^{\prime}}\hbox{-}x_{4}.$ Then $C$ is a hole on more than seven vertices and $W=(C,h_{4})$ is a special wheel in $G$ where $Z(W)=\{d_{i^{\prime}},d_{j^{\prime}},h_{1},h_{4},h_{7},x_{4}\}$; in particular, $Z(W)$ is a $6$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$. By Theorem 3.2, $N[Z(W)]$ separates $d_{i}$ and $d_{j}$.
Finally, assume that $(H,X)$ is a consistent alignment. Recall that $(D_{1},X)$ is also a consistent alignment, and that $(D_{1},X)$ and $(H,X)$ give the same order $x_{1},\ldots,x_{7}$ on $X$. Let $R$ be the unique path in $G$ from $x_{1}$ to $x_{7}$ with $R^{*}\subseteq H$. Then $C=d_{i}\hbox{-}x_{1}\hbox{-}R\hbox{-}x_{7}\hbox{-}d_{j}\hbox{-}D_{1}\hbox{-}d_{i}$ is a hole on more than seven vertices in $G$. Also, since $G$ is (theta, prism)-free, it follows that either one of $(D_{1},X)$ and $(H,X)$ is spiky and the other is triangular, or at least one of $(D_{1},X)$ and $(H,X)$ is wide. In the former case, $W=(C,x_{4})$ is a special wheel (see Figure 14). It follows from Theorem 3.2 that $Z(W)$ is a $6$-subset of $D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$ such that $N[Z(W)]$ separates $d_{i}$ and $d_{j}$. In the latter case, $W=(C,x_{4})$ is a non-special wheel (see Figure 15). Since $G$ is $K_{1,t}$-free, it follows that $Z(W)=N_{C}[x_{4}]\subseteq D_{2}\cup\{d_{k}:i+2\leq k\leq j-2\}\cup\{x_{4}\}$ has cardinality at most $2t$. Moreover, by Theorem 3.1, $N[Z(W)]$ separates $d_{i}$ and $d_{j}$. This completes the proof of Theorem 7.1. ∎
We also need the following result from [tw15]:
Theorem 7.2 (Chudnovsky, Gartland, Hajebi, Lokshtanov, Spirkl [tw15]).
[…]
  • $N[Y]$ is a $w$-balanced separator in $G$.
[…]
8. Acknowledgement
Part of this work was done when Nicolas Trotignon visited Maria Chudnovsky at Princeton University with the generous support of the H2020-MSCA-RISE project CoSP- GA No. 823748.